OpenStudy (anonymous):

The measures of each exterior angle in a regular hexagon is (4x + 15)°. Find the value of x.

OpenStudy (anonymous):

You know, given a polygon of n sides, the total interior angle measure in degrees sums up to 180 ( n - 2 ) Since this is a hexagon, n=6, so the total angle measure is 180 ( 6 - 2 ) = 180 ( 4 ) = 720 Since this is a regular hexagon, each interior angle has the same measure, and so, each angle should measure 720 / 6 = 120 Then 4x + 15 = 120 Now solve for x :)

OpenStudy (anonymous):

thanx peter pna

OpenStudy (anonymous):

No problem :)